Explain how to use FCF evaluation model to find the price per share of common equity.
Answer #1

FCF model is used to find the value of business by discounting the future cash flows of the business. Thus the FCF next year is discounted by the required rate of return to arrive at the total business value. The value of debt is subtracted from the Business value to get value of equity. This value is then divided by the number of shares to arrive at the price per share.

Example: The total business value= FCF1/ (WACC-growth)

=$100/(10%-6%)

= $2500

If the value of debt is $500, value of equity= 2500- 500 = $2000

Value of equity divided by the number of shares will give the price per share.